OpenStudy (anonymous):

Please help :X Find theta in radians if 0 ≤ theta ≤ 2 pi and tan square theta = 1/3

mathslover (mathslover):

May be some one help you :soon \[\huge{\mathbb{Welcome}\textbf{To}\mathbb{Open}\textbf{Study}}\] Also dont forget to give the user who answers your question correctly by clicking on best response !!! thanks!

OpenStudy (campbell_st):

if \[\tan ^2 (\theta) = \frac{1}{3}\] then \[\tan(\theta) = \pm \frac{1}{\sqrt{3}}\]

OpenStudy (anonymous):

thank u :)

OpenStudy (campbell_st):

the value of \[\tan^{-1}(\frac{1}{\sqrt{3}}) = \frac{\pi}{6}\] so there is an angle in every quadrant to find them just complete the calculations below so first quadrant the answer is above 2nd quadrant its \[\pi - \frac{\pi}{6}\] 3rd quadrant \[\pi + \frac{\pi}{6}\] 4th quadrant \[2\pi - \frac{\pi}{6}\]